OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 95 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 90? (Hint: The average of a list of numbers is their sum divided by the number of numbers in the list.)

OpenStudy (anonymous):

\[\frac{85+88+87+95+x}{5} >= 90\]\[\frac{355+x}{5} >=\ 90\]5*90=<355x x>=450/355 \[x \ge 1.27\]

OpenStudy (anonymous):

Solve the following equation for x:\[\frac{85+88+87+95+x}{5}=90 \]x = 95\[\frac{85+88+87+95+95}{5}=90 \]